OpenStudy (anonymous):

and this one too. 2a+4/3 - a-1/4 = 6

OpenStudy (anonymous):

a+59/12

OpenStudy (anonymous):

a=59/12 *

OpenStudy (anonymous):

but the other math solvers got 59/36 as their answer.

OpenStudy (anonymous):

why are your answers different? thanks.

OpenStudy (anonymous):

2a-a+4/3-1/4=a+13/12 a+13/12=6 ===> a=6-13/12=59/12

OpenStudy (anonymous):

okay thanks